# Language-Guided Self-Supervised Video Summarization Using Text Semantic Matching Considering the Diversity of the Video

### v1

@@ Abstract.
Current video summarization methods primarily depend on supervised computer vision techniques, which demands time-consuming manual annotations. Further, the annotations are always subjective which make this task more challenging. To address these issues, we analyzed the feasibility in transforming the video summarization into a text summary task and leverage Large Language Models (LLMs) to boost video summarization. This paper proposes a novel self-supervised framework for video summarization guided by LLMs. Our method begins by generating captions for video frames, which are then synthesized into text summaries by LLMs. Subsequently, we measure semantic distance between the frame captions and the text summary. It’s worth noting that we propose a novel loss function to optimize our model according to the diversity of the video. Finally, the summarized video can be generated by selecting the frames whose captions are similar with the text summary. Our model achieves competitive results against other state-of-the-art methods and paves a novel pathway in video summarization.
@@ 1. Introduction
Figure 1. The overview of our proposed framework. We first generate captions from individual video frames using a pre-trained image captioning model, Generative Image-to-text Transformer (GIT) (Wang et al., 2022). Then the text summary is created by Generative Pre-trained Transformer 4 (GPT-4) (OpenAI, 2023). Based on the captions, diversity of the input video is calculated. Using the optimal loss function according to the diversity of the video, our model calculates semantic distance between individual captions and the text summary.
Video summarization involves distilling a full-length video into a concise version that encapsulates the most crucial or engaging elements of the original. The goal is to produce a summary that is brief yet delivers a cohesive grasp of the principal themes or narratives of the video. Video summarization has emerged as an important research topic in today’s fast-paced information society for two main reasons: 1) There has been an unprecedented increase in video content across social media platforms. This includes not only professional productions such as news broadcasts, live concerts, and sports events but also user-generated content. Dominant platforms especially YouTube and Instagram have become integral to various facets of our daily lives, and they are expected to maintain their far-reaching impact [^1] https://www.statista.com/statistics/1061017/digital-video-viewers-number-worldwide/ Accessed 2024-1-20. 2) The overwhelming volume of available video content, coupled with the modern demand for rapid assimilation of extensive information, underscores the growing necessity for video summarization technology. As it becoming increasingly time-consuming for individuals to consume and process all the available material, video summarization is proving to be an essential development within multimedia and computer vision to address these societal demands.
@@
However, video summarization is inherently complex due to the diverse content and subjectivity involved in identifying key segments within extensive footage. Sophisticated analytical approaches are required to discern these crucial shots. With the evolution of deep neural network architectures, the accuracy of computer vision-based approach has significantly improved. Techniques such as Convolutional Neural Networks (CNNs) for image analysis (Zhou et al., 2018; Mahasseni et al., 2017; Jung et al., 2020; Ghauri et al., 2021), Recurrent Neural Networks (RNNs) for temporal sequence modeling (Zhang et al., 2016; Zhao et al., 2021; Zhong et al., 2022), and the attention mechanisms for highlighting important features (Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023) have enabled effective summarization models. Numerous studies have pursued supervised learning approaches in video summarization, seeking models of heightened accuracy (He et al., 2023; Li et al., 2023; Ghauri et al., 2021; Narasimhan et al., 2021). Nonetheless, such supervised techniques are dependent on extensive annotated datasets, the creation of which can be labor-intensive and intrinsically subjective.
@@
Conversely, the latest progress in Large Language Models (LLMs), including the Generative Pre-trained Transformer 4 (GPT-4) (OpenAI, 2023) and Large Language Model Meta AI 2 (LLaMA 2) (Touvron et al., 2023), has greatly advanced text summarization, enabling the generation of accurate summaries in zero-shot scenarios. This breakthrough in LLMs has opened up new possibilities to video summarization, offering potential solutions to the challenges associated with supervised methods. Furthermore, due to the advancements in Vision-and-Language models, integrating vision and language has become more straightforward. This progress has simplified tasks such as Visual Question Answering (VQA) (Wang et al., 2023), image-to-text (Wang et al., 2022) or text-to-image task (Rombach et al., 2022). Image captioning models, which produce descriptive captions for images, represent a captivating convergence of computer vision and natural language processing (NLP), and are instrumental for tasks that translate visual content into text.
However, leveraging these advanced capabilities for video summarization remains underexplored. In this paper, we delve into the potential of LLMs for video summarization, an approach that has shown promise in various unsupervised scenarios (Yoo et al., 2021; Meng et al., 2022). We develop novel methods to effectively harness LLMs for video summarization. Our proposed method follows a completely different framework from conventional video summarization methods which mainly rely on video frames to calculate frame-level importance scores (frame-level scores). The overview of our proposed method is shown in Figure 1. Our process begins by generating descriptive captions for individual video frames using a pre-trained image captioning model. Then, these generated captions are fed into the LLMs in chronological order, and the resulting coherent text summary is used as reference data. We propose a self-supervised training approach that enables an encoder to calculate semantic distance between the frame captions and the overall text summary within a semantic embedding space. During this phase, the model is optimized using the proposed loss function, tailored to the video’s diversity, hence fostering the development of a robust video summarization model. We termed this proposed loss function the Preserving Diversity Loss (PDL).
Briefly, our contributions are as follows:
  • We are the first to explore the feasibility of transforming a video summarization task into an NLP task. In this context, We propose a novel NLP-based self-supervised video summarization method that leverages the advantages in NLP for video summarization.
  • We propose a novel PDL function that considers the diversity of videos for video summarization.
  • Experimental results on SumMe and TVSum datasets confirm the effectiveness of our proposed model.
@@ 2.1. Video Summarization
Video summarization involves processing a sequence of video frames and producing a binary vector that indicates which shots to be included in the summary. Traditional video summarization methods typically rely on neural networks and can be categorized into supervised and unsupervised approaches.
Supervised learning for video summarization involves methods that use large-scale, manually annotated frame-level data for training (He et al., 2023; Li et al., 2023; Ghauri et al., 2021). MSVA (Ghauri et al., 2021) extracts various visual features from frame images, such as static and dynamic features. A2Summ (He et al., 2023) aligns and attends the multimodal input by an alignment-guided self-attention module to make the use of cross-modal correlation. In SSPVS (Li et al., 2023), self-supervised learning is conducted on both the text encoder and the image encoder during pre-training, with the training data being used for multi-stage fine-tuning in downstream tasks. However, constructing datasets with manual annotations not only requires significant time and effort but also necessitates measures to reduce subjectivity among multiple annotators. Therefore, creating training data is challenging in the field of video summarization (Otani et al., 2019).
In unsupervised learning for video summarization (Zhou et al., 2018; Mahasseni et al., 2017; Jung et al., 2020; Zhang et al., 2016; Zhao et al., 2021; Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023), models calculate frame-level scores using frame images without the need for annotated data. Recent unsupervised video summarization methods employ Generative Adversarial Network (GAN) based methods (Mahasseni et al., 2017; Jung et al., 2020, 2019), RNNs based methods (Zhang et al., 2016; Zhao et al., 2021; Zhong et al., 2022) and self-attention module based methods (Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023; Jung et al., 2019) to calculate frame-level scores. SAM-GAN (Mahasseni et al., 2017) uses GAN to select a subset of keyframes, aiming to generate summaries that closely resemble the original video. DSAVS (Zhong et al., 2022) calculates the similarity between the caption and frame images within the same semantic space, employing Long Short Term Memory (LSTM) module and a self-attention module. RSSUM (Abbasi and Saeedi, 2023) adopts self-supervised learning by training an encoder to reconstruct missing video sections using rule-based masked operations. However, there is still a gap between achieved accuracy and real-world applications. Therefore, more researches are needed.
@@ 2.2. Text Summarization
Text summarization can be divided into two types: extractive summarization and abstractive summarization. Extractive summarization entails selecting significant sentences or phrases from a document and combining them to create a text summary. MatchSum (Zhong et al., 2020) formulates this as a semantic text matching problem, where the correct summary is semantically embedded closer to the original document compared to other candidate summaries. Specifically, the model employs a Siamese-BERT architecture that is based on the Siamese network (Bromley et al., 1994). The Siamese-BERT consists of two Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) with shared-weights and a cosine similarity layer.
Abstractive summarization involves concisely paraphrasing the main content from the sentences within the input document while retaining the important parts. Recent advancements have been marked by the development of LLMs such as GPT-4 (OpenAI, 2023) and LLaMA 2 (Touvron et al., 2023), which are built upon the Transformer (Vaswani et al., 2017) architecture and have been pre-trained on a vast amount of datasets. These models possess advanced language understanding and generation capabilities, improving the ability to accurately summarize long documents. Also, it is reported that these performances are equivalent to human-written summaries (Zhao et al., 2023; Tam et al., 2023). As a result, text summarization technology has become capable of accurately summarizing more complex and lengthy texts in zero-shot settings.
@@ 3.1. Framework of our method
As previously mentioned, our motivation is to leverage the full potential of LLMs in multi-modal learning. To this end, we propose a novel method to transform video summarization task into an NLP task, which enables representing videos linguistically. Figure 1 provides the overview of our framework. Our method begins by generating descriptive captions for downsampled individual video frames by using a pre-trained image captioning model, Generative Image-to-text Transformer (GIT) (Wang et al., 2022). Then, these captions are synthesized into a coherent text summary by using GPT-4 (OpenAI, 2023). Afterwards, the number of downsampled captions is denoted as $n$, with $C_{i}$ representing the frame captions for $i\in\{1,2,\dots,n\}$, and $T$ representing the generated text summary. We begin with encoding each of the $n$ frame captions and the generated text summary using pre-trained Sentence-BERT (Reimers and Gurevych, 2019). The encoded captions are denoted as $E_{C_{i}}$, and the encoded text summary as $E_{T}$. Then, we perform deep metric learning on the Siamese-Sentence-BERT architecture (Reimers and Gurevych, 2019; Bromley et al., 1994), which consists of two pre-trained Sentence-BERTs with shared-weights and a cosine similarity layer. After the model is trained, the model inputs the pair of individual captions, $C_{i}$ and the text summary, $T$. The encoded captions are denoted as $E^{\prime}_{C_{i}}$ and the encoded text summary is denoted as $E^{\prime}_{T}$. The similarity score $s_{i}$ between $E^{\prime}_{C_{i}}$ and $E^{\prime}_{T}$ is calculated using the cosine similarity, calculated by Eq. (1) as follows:
@@
3.2. Proposed PDL
Overall PDL. We introduce a novel PDL, denoted as $\mathcal{L}_{\rm PDL}$, tailored for accommodating the diversity of the video into video summarization. The PDL mainly contains two components: a margin ranking loss ${L}_{m}$ and a regularization loss ${L}_{s}$ (Mahasseni et al., 2017). Mathematically, $\mathcal{L}_{\rm PDL}$ is calculated as follows:
(3) $\mathcal{L}_{\rm PDL}=\mathcal{L}_{m}+\lambda\mathcal{L}_{s},$
where $\lambda$ is a value to balance the two terms, which is decided by the video diversity.
Specifically, the frame is first represented linguistically by an image captioning model. Then, we calculate the semantic similarity of frame captions using ${L}_{m}$. We also apply a sparsity regularization loss, ${L}_{s}$ to constrain the number of keyframes. Finally, we introduce a unique diversity metric, denoted as $\lambda$, to dynamically adjust the balance between two losses. When the diversity of the video is high, incorporating the regularization loss becomes unnecessary, as the inherent diversity ensures a rich and comprehensive representation of the video content. Conversely, if similar captions are prevalent, making it challenging to differentiate based on similarity, it becomes necessary to reduce the contribution of the margin ranking loss. Details of the margin ranking loss and regularization loss are given below.
Margin Ranking Loss. Margin ranking loss is adopted to make the semantically close captions closer, and increase the score difference between text summary and semantically different captions, which is defined as:
(4) $\mathcal{L}_{m}=\frac{1}{n}\sum_{i=1}^{n}\max(0,-y_{i}\times(S_{i}-S_{\rm{avg}})+m),$
where $m$ represents the minimum score difference the model should maintain between the positive and negative classes. In other words, the larger the $m$ is, the greater the difference in scores between the positive and negative classes when using the model after training would become. The $S_{\rm{avg}}$ is the average value of $S_{i}$ and is calculated as follows:
@@ 3.2. Proposed PDL
Also, $y_{i}$ is a target label as shown in Eq. (6). We classify $S_{i}$ into two classes according to the following rules. It is considered a score belonging to the positive class if $S_{i}$ is greater than $S_{\rm{avg}}$, and conversely, it is considered a score belonging to the negative class.
(6) $y_{i}=\begin{cases}1&\text{if }S_{i}\geq S_{\rm{avg}}\\ -1&\text{if }S_{i}<S_{\rm{avg}}.\end{cases}$
Regularization Loss. We selectively apply this sparsity regularization loss to ensure a more concise and informative summary as exiting works (Zhong et al., 2022; Mahasseni et al., 2017; Apostolidis et al., 2022), which is defined as:
@@
where $\epsilon$ is a hyperparameter that specifies the proportion of frames to be selected as keyframes and set as 0.3 as existing works (Zhong et al., 2022; Mahasseni et al., 2017). By constraining the number of key frames, summaries become more concise and relevant, avoiding redundancy and focusing on the most informative parts of the video.
Value of $\lambda$.

### v2

@@ Abstract.
Current video summarization methods rely heavily on supervised computer vision techniques, which demands time-consuming and subjective manual annotations. To overcome these limitations, we investigated self-supervised video summarization. Inspired by the success of Large Language Models (LLMs), we explored the feasibility in transforming the video summarization task into a Natural Language Processing (NLP) task. By leveraging the advantages of LLMs in context understanding, we aim to enhance the effectiveness of self-supervised video summarization. Our method begins by generating captions for individual video frames, which are then synthesized into text summaries by LLMs. Subsequently, we measure semantic distance between the captions and the text summary. Notably, we propose a novel loss function to optimize our model according to the diversity of the video. Finally, the summarized video can be generated by selecting the frames with captions similar to the text summary. Our method achieves state-of-the-art performance on the SumMe dataset in rank correlation coefficients. In addition, our method has a novel feature of being able to achieve personalized summarization.
@@ 1. Introduction
Figure 1. Overview of our proposed framework. We take only videos as input and first generate captions from individual frames using a pre-trained image captioning model, the Generative Image-to-text Transformer (GIT) (Wang et al., 2022). The text summary is then created by GPT-4 (OpenAI, 2023). The semantic distance between individual captions and the text summary is calculated using the proposed Preserving Diversity Loss (PDL) to calculate frame-level scores. Finally, the frame-level scores are aggregated into scene-level scores, and the knapsack problem is solved to select a subset of scenes, thereby creating the video summary.
Video summarization involves distilling a full-length video into a concise version that encapsulates the most crucial or engaging elements of the original. The goal is to produce a summary that is brief yet delivers a cohesive grasp of the principal themes or narratives of the video. Video summarization has emerged as an important research topic in today’s fast-paced information society for two main reasons: 1) There has been an unprecedented increase in video content across social media platforms. This includes not only professional productions such as news broadcasts, live concerts, and sports events but also user-generated content. Dominant platforms especially YouTube and Instagram have become integral to various facets of our daily lives, and they are expected to maintain their far-reaching impact [^1] https://www.statista.com/statistics/1061017/digital-video-viewers-number-worldwide/. 2) The overwhelming volume of available video content, coupled with the modern demand for rapid assimilation of extensive information, underscores the growing necessity for video summarization technology. As it becoming increasingly time-consuming for individuals to consume and process all the available material, video summarization is proving to be an essential development within multimedia and computer vision to address these societal demands.
@@
However, video summarization is inherently complex due to the diverse content and subjectivity involved in identifying key segments within extensive footage. Sophisticated analytical approaches are required to discern these crucial shots. With the evolution of deep neural network architectures, the accuracy of computer vision-based approach has significantly improved. Techniques such as Convolutional Neural Networks (CNNs) for image analysis (Zhou et al., 2018; Mahasseni et al., 2017; Jung et al., 2020; Ghauri et al., 2021), Recurrent Neural Networks (RNNs) for temporal sequence modeling (Zhang et al., 2016; Zhao et al., 2021; Zhong et al., 2022), and the attention mechanisms for highlighting important features (Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023) have enabled effective summarization models. Numerous studies have pursued supervised learning approaches in video summarization, seeking models of heightened accuracy (He et al., 2023; Li et al., 2023a; Ghauri et al., 2021; Argaw et al., 2024; Tian et al., 2024; Narasimhan et al., 2021; Otani et al., 2017; Jiang and Mu, 2022; Son et al., 2024). However, these supervised video summarization methods rely heavily on large amounts of human-generated annotated data, which is time-consuming and subjective. As a result, low-quality labeled data can significantly restrict the performance.
@@
Conversely, the latest progress in Large Language Models (LLMs), including the Generative Pre-trained Transformer 4 (GPT-4) (OpenAI, 2023) and Large Language Model Meta AI 2 (LLaMA 2) (Touvron et al., 2023), has greatly advanced text summarization, enabling the generation of accurate summaries in zero-shot scenarios. This breakthrough in LLMs has opened up new possibilities to video summarization, offering potential solutions to the challenges associated with supervised methods. Furthermore, due to the advancements in Vision-and-Language models, integrating vision and language has become more straightforward. This progress has simplified tasks such as Visual Question Answering (VQA) (Wang et al., 2023a), image-to-text (Wang et al., 2022) or text-to-image task (Rombach et al., 2022). Image captioning models, which produce descriptive captions for images, represent a captivating convergence of computer vision and natural language processing (NLP), and are instrumental for tasks that translate visual content into text.
To address the aforementioned issues in the video summarization task, this paper explores the potential of LLMs in video summarization, which have demonstrated effectiveness in natural language processing and contextual understanding (Yoo et al., 2021; Meng et al., 2022; Wang et al., 2023b; OpenAI, 2023; Touvron et al., 2023; Zhao et al., 2023a). To this end, we propose a novel self-supervised method that effectively leverages LLMs for video summarization and introduce a loss function tailored to the video’s diversity, thereby fostering the development of a robust video summarization model. We term this proposed loss function the Preserving Diversity Loss (PDL). Briefly, our contributions are as follows:
  • We are the first to transform a video summarization task into a semantic textual similarity task in a self-supervised way.
  • We propose a novel self-supervised video summarization method that leverages LLMs, unlocking their potential in natural language and contextual/semantic understanding to enhance video summarization.
  • We mathematically analyze the characteristics of the dataset and adjust the balancing parameter based on the diversity of the video to further enhance our model. Experimental results verified the effectiveness of our proposed model.
  • Our proposed framework is versatile and can incorporate user specifications to generate user-guided video summaries, which has been very difficult in previous approaches.
@@ 2.1. Video Summarization
Video summarization involves processing a sequence of video frames and producing a binary vector that indicates which shots to be included in the summary. Conventional video summarization methods can be categorized into unsupervised (Zhou et al., 2018; Mahasseni et al., 2017; Jung et al., 2020; Zhang et al., 2016; Zhao et al., 2021; Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023; Wang et al., 2024) and supervised learning methods (He et al., 2023; Li et al., 2023a; Ghauri et al., 2021; Argaw et al., 2024; Tian et al., 2024; Narasimhan et al., 2021; Otani et al., 2017; Jiang and Mu, 2022; Son et al., 2024). Unsupervised video summarization methods calculate frame-level scores by using frame images without relying on annotated data. Recent unsupervised video summarization methods employ Generative Adversarial Network (GAN) based methods (Mahasseni et al., 2017; Jung et al., 2020, 2019), RNNs based methods (Zhang et al., 2016; Zhao et al., 2021; Zhong et al., 2022) and self-attention module based methods (Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023; Jung et al., 2019) to calculate frame-level scores. SAM-GAN (Mahasseni et al., 2017) uses GAN to select a subset of keyframes, aiming to generate summaries that closely resemble the original video. DSAVS (Zhong et al., 2022) calculates the similarity between the caption and frame images within the same semantic space, employing Long Short Term Memory (LSTM) module and a self-attention module. RSSUM (Abbasi and Saeedi, 2023) adopts self-supervised learning by training an encoder to reconstruct missing video sections using rule-based masked operations.
Supervised learning for video summarization involves methods that use large-scale, manually annotated frame-level data for training (He et al., 2023; Li et al., 2023a; Ghauri et al., 2021; Argaw et al., 2024; Tian et al., 2024; Narasimhan et al., 2021; Otani et al., 2017; Jiang and Mu, 2022; Son et al., 2024). MSVA (Ghauri et al., 2021) extracts various visual features from frame images, such as static and dynamic features. A2Summ (He et al., 2023) aligns and attends the multimodal input by an alignment-guided self-attention module to make the use of cross-modal correlation. In SSPVS (Li et al., 2023a), self-supervised learning is conducted on both the text encoder and the image encoder during pre-training, with the training data being used for multi-stage fine-tuning in downstream tasks. However, constructing manually annotated datasets is not only time-consuming and expensive, but also challenging. This process necessitates reducing subjectivity among annotators. Therefore, creating training data for video summarization is not a long-term solution (Otani et al., 2019; Sul et al., 2023; Terbouche et al., 2023). It is essential to develop self-supervised video summarization technologies for real-world applications.
@@ 2.2. Text Summarization
Text summarization can be divided into two types: extractive (Zhong et al., 2020; Xu et al., 2022; Sosea et al., 2023; Liu and Lapata, 2019) and abstractive summarization (Touvron et al., 2023; OpenAI, 2023; Team et al., 2023; Lewis et al., 2020). Extractive summarization entails selecting significant sentences or phrases from a document and combining them to create a text summary. MatchSum (Zhong et al., 2020) formulates this as a semantic text matching problem, where the correct summary is semantically embedded closer to the original document compared to other candidate summaries. Specifically, the model employs a Siamese-BERT architecture that is based on the Siamese network (Bromley et al., 1994). The Siamese-BERT consists of two Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) with shared-weights and a cosine similarity layer.
Abstractive summarization involves concisely paraphrasing the main content from the sentences within the input document while retaining the important parts. Recent advancements have been marked by the development of LLMs such as GPT-4 (OpenAI, 2023) and LLaMA 2 (Touvron et al., 2023), which are built upon the Transformer (Vaswani et al., 2017) architecture and have been pre-trained on a vast amount of datasets. These models possess advanced language understanding and generation capabilities, improving the ability to accurately summarize long documents. Also, it is reported that these performances are equivalent to human-written summaries (Zhao et al., 2023b; Tam et al., 2023). As a result, text summarization technology has become capable of accurately summarizing more complex and lengthy texts in zero-shot settings.
@@ 3.1. Framework of our method
As mentioned before, our motivation is to leverage LLMs for unsupervised video summarization, freeing the process from the burden of extensive data annotation and the subjective errors associated with it. We transform the video summarization task into an NLP task, enabling videos to be represented linguistically and fully taking advantage of LLMs in contextual/semantic understanding. Figure 1 provides the overview of our framework. Our method begins by generating descriptive captions for downsampled individual video frames by using a pre-trained image captioning model, Generative Image-to-text Transformer (GIT) (Wang et al., 2022). Then, these captions are synthesized into a coherent text summary by using GPT-4 (OpenAI, 2023). Afterwards, the number of downsampled captions is denoted as $n$, with $C_{i}$ representing the frame captions for $i\in\{1,\dots,n\}$, and $T$ representing the generated text summary. We begin with encoding each of the $n$ frame captions and the generated text summary using pre-trained Sentence-BERT (Reimers and Gurevych, 2019) as shown in Figure 2. The encoded captions are denoted as $E_{C_{i}}$, and the encoded text summary as $E_{T}$. Then, we perform deep metric learning on the Siamese-Sentence-BERT architecture (Reimers and Gurevych, 2019; Bromley et al., 1994), which consists of two pre-trained Sentence-BERTs with shared-weights and a cosine similarity layer. After the model is trained, the model inputs the pair of individual captions, $C_{i}$ and the text summary, $T$. The encoded captions are denoted as $E^{\prime}_{C_{i}}$ and the encoded text summary is denoted as $E^{\prime}_{T}$. The similarity score $s_{i}$ between $E^{\prime}_{C_{i}}$ and $E^{\prime}_{T}$ is calculated using the cosine similarity, calculated by Eq. (1) as follows:
@@
In terms of computational complexity, we calculate the semantic textual similarity for individual captions and the text summary separately, resulting in a complexity of $\mathcal{O}(d\cdot n)$, where $d$ is the hidden dimension.
3.2. Loss function
The loss function of our model consists of two components, an improved margin ranking loss and a sparsity regularization loss (Mahasseni et al., 2017)). Details of these two components will be given below.
Improved Margin Ranking Loss. The conventional margin ranking loss is used for learning the relative relevance between different items. Normally, it has positive and negative pairs as input. In our task, we only care about the absolute relevance. Therefore, we propose an improved margin ranking loss tailored for the video summarization task. It aims to increase the score difference between captions and the text summary, which is defined as:
(3) $\mathcal{L}_{m}=\frac{1}{n}\sum_{i=1}^{n}\max(0,-\lvert S_{i}-S_{\rm{avg}}\rvert+m),$
where $S_{\rm{avg}}$ is the average value of $S_{i}$ and is calculated as follows:
@@ 3.2. Loss function
Also, $m$ is the minimum score difference the model should maintain between the positive and negative classes. We classify $S_{i}$ into two classes according to the following rules. It is considered a score belonging to the positive class if $S_{i}$ is greater than $S_{\rm{avg}}$, and conversely, it is considered a score belonging to the negative class. Therefore, the larger the $m$ is, the greater the difference in scores between the positive and negative classes when using the model after training would become.
Regularization Loss. We apply this sparsity regularization loss to ensure a more concise and informative summary as exiting works (Apostolidis et al., 2022; Zhong et al., 2022; Zhou et al., 2018; Mahasseni et al., 2017; Wei et al., 2018), which is defined as:
@@
where $\epsilon$ is a hyperparameter that specifies the proportion of frames to be selected as key frames and set to 0.3 as in existing works (Zhong et al., 2022; Mahasseni et al., 2017; Wei et al., 2018). By constraining the number of key frames, summaries become more concise and relevant, avoiding redundancy and focusing on the most informative parts of the video.
3.3. Impact of the loss functions
Figure 3. Effect of the fixed contribution value ($\alpha$) of the regularization loss (Mahasseni et al., 2017). A higher value of Spearman’s $\rho$ indicates a better video summary.